### v2

# Adaptive Hybrid Model Pruning in Federated Learning through Loss Exploration

Abstract
The rapid proliferation of smart devices coupled with the advent of 6G networks has profoundly reshaped the domain of collaborative machine learning. Alongside growing privacy-security concerns in sensitive fields, these developments have positioned federated learning (FL) as a pivotal technology for decentralized model training. Despite its vast potential, specially in the age of complex foundation models, FL encounters challenges such as elevated communication costs, computational constraints, and the complexities of non-IID data distributions. We introduce AutoFLIP, an innovative approach that utilizes a federated loss exploration phase to drive adaptive hybrid pruning, operating in a structured and unstructured way. This innovative mechanism automatically identifies and prunes model substructure by distilling knowledge on model gradients behavior across different non-IID client losses topology, thereby optimizing computational efficiency and enhancing model performance on resource-constrained scenarios. Extensive experiments on various datasets and FL tasks reveal that AutoFLIP not only efficiently accelerates global convergence, but also achieves superior accuracy and robustness compared to traditional methods. On average, AutoFLIP reduces computational overhead by 48.8% and communication costs by 35.5%, while improving global accuracy. By significantly reducing these overheads, AutoFLIP offer the way for efficient FL deployment in real-world applications for a scalable and broad applicability.
1 Introduction
The proliferation of smart devices at the network edge, coupled with advancements in 6G networks, has created a decentralized setting [31, 47]. Multiple participants store their data locally, which offers an opportunity for collaborative model training, enhancing robustness and generalization. Distributing the computational load across these devices results in faster training times and lower energy consumption compared to centralized approaches [56, 3]. However, collaborative Machine Learning (ML) faces significant challenges [8]. Efficient communication and coordination among participants are crucial, as each device holds only a subset of the data. This requires designing algorithms that minimize data exchange while ensuring high-quality model convergence. Device heterogeneity, including differences in computational power, storage, and bandwidth, further complicates distributed training. Algorithms must adapt to such environments to scale up distributed learning. Privacy and security concerns, along with regulations like the European GDPR [1], the EU AI Act [9] and the U.S. Secure AI [54] Act add another layer of complexity [21]. With sensitive data distributed across various devices, ensuring the privacy of individual data points becomes essential [14]. For example, medical data stored in hospitals and personal devices is valuable for training diagnostic models but is also subject to strict privacy and security regulations [46].
In this context, Federated Learning (FL) [65] emerges as an effective strategy for training always more complex DL models while preserving the privacy of the data. FL facilitates collaborative model training across multiple devices without exposing local data. A central server, i.e., a global model, coordinates this process by aggregating updates from locally trained models, which ensures a secure learning environment. Current FL research focuses on enhancing privacy and adapting ML workflows for specific uses, often with predetermined ML model configurations. Tasks related to computer vision may involve well-known neural network (NN) architectures like VGG-16 [49] (138 million parameters) or ResNet-50 [19] (25.6 million parameters). However, these complex NN risk overfitting, especially with small training data sizes.
In the era of foundation models [3] becoming the norm in machine learning development, FL systems typically expect clients to have high-speed processors and sufficient computational power for local calculations and parameter updates. Yet, many edge devices, such as smartphones, wearable, and sensors, have limited computing and memory capacities, posing a challenge to DL model training systems [21]. Additionally, communicating DL models with millions of parameters presents significant obstacles for FL transmission [48, 2]. Therefore, using FL effectively with edge devices that have limited computational capabilities, while maintaining efficient communication, remains an active research question. FL’s effectiveness is further hindered by the prevalence of non-IID data in real-world scenarios [65, 23, 27]. non-IID data refers to the unique statistical properties of each client’s dataset, reflecting their varied environments. This creates conflicting training goals for local and global models, leading to convergence towards different local optima. As a result, client model updates become biased, impeding global convergence [23]. These challenges underscore the need for personalized and innovative approaches in FL, particularly in optimizing and compressing models to improve inference time, communication cost, energy efficiency, and complexity, all while maintaining satisfactory accuracy.
Our contribution. We introduce a novel automated federated learning approach via adaptive hybrid pruning (AutoFLIP), which uses a novel loss exploration mechanism to automatically prune and compress DL models. In our assumed single-server architecture, each client operates on the same initial deep NN structure that automatically prunes itself at each round, based on the extraction of shared knowledge from the federated loss exploration for an informed model compression. Specifically, by analyzing the variability of gradients during a preliminary local loss exploration phase, which provides insights into gradient behaviors on the loss landscapes across clients, and subsequent information aggregation, the DL models involved in a FL round are hybridly pruned automatically. This strategy allows for dynamically reducing the complexity of the models in FL environments, thereby optimizing performance with limited computational resources at the client level. With our experiments over various datasets, tasks, and realistic non-IID scenarios, we provide strong evidence of the effectiveness and efficiency of AutoFLIP.
Reproducibility. Our code for reproducing the experiments is available on Anonymous GitHub.111 https://github.com/ChristianInterno/AutoFLIP
2 Background and Related Work
Pruning in Deep Learning. Following the assumption that a DL model can contain a sub-network that represents the performance of the entire model after being trained, model pruning is a good strategy to reduce computational requirements of resource-constrained devices [41, 29, 25]. Most pruning approaches balance accuracy and sparsity during the inference stage by calculating the importance scores of parameters in a well-trained NN and removing those with lower scores. These scores can be derived from weight magnitudes [25, 17], first-order Taylor expansion of the loss function [42, 39], second-order Taylor expansion [29, 18, 40], and other variants [37, 50].
Another recent research direction in NN pruning focuses on improving training efficiency, divided into two categories: pruning at initialization and dynamic sparse training. Pruning at initialization involves pruning the original full-size model before training based on connection sensitivity [30], Hessian-gradient product [55], and synaptic flow [51]. However, since this method does not involve training data, the pruned model may be biased and not specialized for the task. Dynamic sparse training iteratively adjusts the pruned model structure during training while maintaining the desired sparsity [7, 10]. This approach requires memory-intensive operations due to the large search space, making it impractical for resource-constrained devices.
Initial attempts to use pruning for deploying deep neural networks on resource-limited devices have utilized pre-trained CNNs in a centralized setting [61, 33]. However, this approach can lead to reduced data privacy, higher costs, poor adaptation to local conditions, suboptimal performance on diverse data, and latency in real-time applications.
Hybrid Pruning in Deep Learning. Hybrid pruning techniques combine structured and unstructured pruning strategies to optimize both performance and efficiency in deep neural networks [36, 44, 13, 15]. Structured pruning [20] removes entire units like neurons, filters, or layers, leading to more hardware-efficient designs that are easier to implement on resource-constrained devices. On the other hand, unstructured pruning [32]focuses on removing individual weights, which can achieve higher sparsity levels and further compress the model, though it may require more complex hardware support [40].
Pruning in Federated Learning. The widely accepted FL standard is known as FedAvg [38]. It distributes a global model to clients for local training and aggregates it by averaging their parameters. Empirical studies have shown the robustness of this approach, even when handling non-convex optimization problems [5]. As a result, it is commonly used as a standard for evaluating newly developed FL protocols. In this study, we will compare the performance of the proposed AutoFLIP method to FedAvg, with different State-of-the-Art (SotA) FL pruning approachs, as tested in [58]. In fact, since data remains locally stored and cannot be shared, traditional centralized pruning approaches that rely on access to training data are not feasible in FL.
In the context of FL, there has been work focused on dynamic active pruning to increase communication efficiency during training. Liu et al. [35], Zhou et al. [64] introduced a method where pruning decisions are made dynamically based on the model’s real-time performance evaluation, which significantly reduces the data exchanged during training but adds computational complexity to client devices. Jiang et al. [26] introduced PruneFL, a FL method that incorporates adaptive and distributed parameter pruning. Their approach utilizes an unstructured method that does not take advantage of the collective insights of participating clients to develop a cooperative structured pruning strategy. This is in contrast to the objectives of AutoFLIP, which seeks to harness client-specific knowledge to facilitate a structured approach to pruning. Lin et al. [34] introduced a novel approach for adaptive per-layer sparsity, however without incorporating any parameter aggregation scheme to reduce the error caused by pruning. This challenge was addressed by Tingting et al. [52] by moving the pruning process to the global model that works on a computationally more powerful server. The pruned model is distributed to each client, where it undergoes training. Subsequently, each client sends back to the server only the updated parameters, restoring the full structure of the model at the server. Although this study includes various parameter selection criteria from the literature, its pruning method does not incorporate the information gathered during model training. This contrasts with our strategy, AutoFLIP, which leverages such information to enhance the pruning process. Yu et al. [62] proposed Resource-aware Federated Foundation Models, focusing on integrating large transformer-based models into FL, with the limitation of not exploring other architectures. Our method, AutoFLIP, diverges by introducing a pruning strategy that avoids the need for continuous evaluation of parameter significance and is universally applicable across various FL aggregation algorithms and model architectures.
3 Preliminaries
Notation: We consider a total number of $C$ clients. At each FL round, $K$ clients are chosen and trained on different batches of size $B$ for $E$ epochs. The total number of rounds is $R$, which represents our termination criterion. For the exploration phase,$C_{\text{exp}}$ is the number of clients selected, which, in this study, we take as the totality $C$ of available clients. The exploration lasts for $E_{\text{exp}}$ epochs.
3.1 Federated Learning
In the conventional FL setting, each client $i$ ($1\leq i\leq K$) possesses its own data distribution $p_{i}(x,y)$, where $x\in\mathbb{R}^{d}$ represents the $d$-dimensional input vector and $y\in\{1,\ldots,M\}$ is the corresponding label from $M$ classes. Each client has a dataset $D_{i}$ with $N_{i}$ data points: $D_{i}=\{(x_{i}^{(1)},y_{i}^{(1)}),\ldots,(x_{i}^{(N_{i})},y_{i}^{(N_{i})})\}$. It is assumed that in a non-IID scenario the data distribution $p_{i}(x,y)$ varies across clients. These data distributions $p_{i}(x,y)$ are sampled from a family $\mathcal{E}$ of distributions. The objective is for the clients to collaboratively train a global model with parameters $W_{\text{global}}$, which will perform predictions on new data. The global loss function for a data point $(x,y)$ is denoted by $\mathcal{L}(W_{\text{global}},x,y)$, where the global objective function to be minimized is defined as: $\mathcal{L}(W_{\text{global}}):=\frac{1}{C}\sum_{i=1}^{C}\mathbb{E}_{(x_{i},y_{i})\sim p_{i}}[\mathcal{L}(W_{\text{global}},x_{i},y_{i})],$ with $\mathbb{E}_{(x_{i},y_{i})\sim p_{i}}$ representing the expected loss over the data distribution $p_{i}$ for each client $i$ with parameters $W$.
The optimization process involves several key steps:
1. Client Selection: A subset of $K$ clients is selected from the total $C$ clients. 2. Local Training: Each selected client $i$ performs local training for $E$ epochs using its local dataset $D_{i}$. The local training aims to minimize the local objective function $\mathcal{L}(W_{i})$ using stochastic gradient descent (SGD): let $W_{i}^{r}$ be the local model parameters of client $i$ at round $r$, the update rule is given by: $W_{i}^{r+1}=W_{i}^{r}-\eta\nabla\mathcal{L}(W_{i}^{r})$, where $\eta$ is the learning rate. 3. Parameter Aggregation: After local training, each client sends its updated parameters $W_{i}^{r+1}$ to the central server. The server aggregates these parameters to form the new global model $W_{global}^{r+1}$ using a weighted average: $W_{global}^{r+1}=\frac{1}{K}\sum_{i=1}^{K}W_{i}^{r+1}$. This iterative process is repeated for $R$ rounds the termination criterion is met.
3.2 Problem Definition and Objective
i) Mitigates noise and biases in the client trajectories: By selectively pruning model parameters based on their contribution to loss topology variability, we aim to the trajectories of the different clients to help converging the global model more effectively despite the heterogeneous data distributions across clients. ii) Enhance Computation and Communication Efficiency: The hybrid pruning serves as a mechanism to decrease the number of model parameters that need to be communicated between clients and the central server. By eliminating less critical parameters and substructure, AutoFLIP reduces communication overhead and accelerates the computational process. This not only expedites the overall FL workflow but also makes it more feasible to deploy in resource-constrained scenarios.
4 Methodology
AutoFLIP is an automatized FL approach that utilizes informed pruning through a federated client loss exploration process. Inspired by the idea of utilizing agents with similar tasks as scouts which explore the conformation of different loss function landscapes from Internò et al. [23], Nikolić et al. [43], AutoFLIP introduces a preliminary step to the FL process, which we term federated loss exploration phase. Here, a $C_{\text{exp}}$ portion of clients (or the totality $C$), which inherit their model structure from the global model, explore for a number of $E_{\text{exp}}$ exploration epochs their loss landscape using its local dataset $D_{i}$. Based on this, for each client $c_{\text{exp}_{i}}$, we compute a local guidance matrix $G_{\text{local}_{i}}$, which records how important a certain parameter $W_{i}$ (weight or bias) is in terms of loss variability 1. Afterward, we aggregate the information collected locally in a global pruning guidance matrix $PG_{\text{global}}$ on the server 2, which will generate an informed pruning mask to guide the pruning of the client models 3. The pruning workflow of AutoFLIP is illustrated in Figure 1. Please note that the initial federated loss exploration, computation of parameter deviations, and definition of local guidance matrices occur only once at the beginning of the FL optimization process as a preliminary procedure. In contrast, the global guidance matrix and subsequent pruning strategy are automatically redefined in each FL round, considering the clients participating in that round.
To summarize, the iterative procedure consists of (1) pruning local models using the updated pruning guidance matrix, (2) training the pruned local models, (3) aggregating the model parameters and (4) evaluating performance and updating the pruning guidance matrix in each FL round.
Figure 1: Illustration of the AutoFLIP pruning procedure. The local guidance matrices are computed a priori through the federated exploration phase. The global guidance matrix is computed by the server by aggregating the elements of the local guidance matrices corresponding to the clients participating in each FL round. The pruning mask is downloaded by the participant clients. All steps preliminary to the FL procedure are denoted in gray, while the steps intrinsic to the FL procedure with pruning are denoted in red.
4.1 Federated Loss Exploration
In AutoFLIP, the model initialization phase is augmented by a crucial federated loss exploration phase, allowing clients to explore their loss function landscapes. We envision each client as an explorer that delves into different regions of their loss landscape. Through this exploration, they can identify crucial dimensions and those that can be disregarded based on their experience by quantifying gradient variability during the exploration. In other words, how much the loss is steep in that direction. Subsequently, they transmit this knowledge to the server, which updates a pruning guidance mask $PG_{\text{global}}$. This knowledge contained in the mask is then distilled among participating clients in each FL round to guide the evolution of client model structures within an informed pruning session.
To construct the mask $PG_{\text{global}}$, we begin with an initial exploration phase conducted on $C_{\text{exp}}$ clients. In this study, we consider $C_{\text{exp}}=C$. In our study, we let explore the clients for $E_{\text{exp}}=150$ epochs. Appendix A provides ablation studies for $C_{\text{exp}}$ and $E_{\text{exp}}$. For each model parameter we evaluate its evolution in the search space during the loss exploration. This evaluation is conducted by calculating the deviation $D_{i,m}$ for the $m^{th}$ parameter of a client model $i$ as the squared difference between the initial ($W_{i,m}^{\text{Initial}}$) and final ($W_{i,m}^{\text{Final}}$) parameter values after $E_{\text{exp}}$ epochs of exploration:
$D_{i,m}=(W_{i,m}^{\text{Initial}}-W_{i,m}^{\text{Final}})^{2}.$ (1)
Using stochastic gradient descent for exploration, the deviation $D_{i,m}$ in Eq. (1) serves as a measure of gradient variability on the loss landscape for parameter $m$ during the preliminary exploration phase before the actual FL procedure. The greater the variation in the parameter space, the faster the improvements in loss: the update rule for a parameter in stochastic gradient descent is $W_{{i,m}}^{(e_{\text{exp}}+1)}=W_{{i,m}}^{(e_{\text{exp}})}-\eta\nabla\mathcal{L}_{i}\left(W_{{i,m}}^{(e_{\text{exp}})};D_{i}\right))$, where $W_{{i,m}}^{(e_{\text{exp}})}$ and $W_{{i,m}}^{(e_{\text{exp}}+1)}$ are the values of the parameter $m$ at the exploration epochs $e_{\text{exp}}$ and $e_{\text{exp}}+1$, $\eta$ is the learning rate, and $\nabla\mathcal{L}_{i}\left(W_{{i,m}}^{(e_{\text{exp}})}\right)$ is the gradient of the loss function of client $i$ with respect to the parameter $m$ at epoch $e_{\text{exp}}$ using its local dataset $D_{i}$. Given the gradient update rule, the deviation in $W_{i,m}$ from the initial to the final exploration epoch $E_{\text{exp}}$ can be approximated to $W_{i,m}^{\text{Final}}-W_{i,m}^{\text{Initial}}\approx-\eta\sum_{t=1}^{E_{\text{exp}}}\nabla\mathcal{L}_{i}\left(W_{i,m}^{(t)};D_{i}\right)$. To ensure non-negativity and highlight larger deviations more severely, we take the square of this value. This squared deviation measure $D_{i,m}$ approximates the square of the sum of gradients affecting the parameter evolution, indicating the significance of parameter updates on loss variability during the exploration phase. By squaring the sum of the gradients, we ensure that the deviation measure is always non-negative and that larger deviations are highlighted more severely than smaller ones.
The $C_{\text{exp}}$ clients compile these deviations into a local matrix $G_{\text{local}}$, whose entries are the deviations for the model parameters. At each FL round, where only $K$ clients are involved, the server aggregates the $G_{\text{local}}$ matrices associated to those client to formulate $G_{\text{global}}$ through a normalization process:
$G_{\text{global}}=\frac{1}{K}\sum_{k=1}^{K}\frac{G_{\text{local}_{k}}-\min(G_{\text{local}})}{\max(G_{\text{local})}-\min(G_{\text{local}})}$ (2)
Here, the minimum and maximum values are taken over all $G_{\text{local}_{k}}$ matrices for $k=1,\ldots,K$. Each element of $G_{\text{global}}$ thus represents the mean normalized deviation for each parameter, scaled between 0 and 1. This process ensures that no single client’s $G_{\text{local}}$ disproportionately influences $G_{\text{global}}$ due the possible presence of outliers in terms of deviations $D_{i,m}$. A value closer to 0 indicates minimal deviation, suggesting gradient stability during the exploration, hence scarce relevance of the parameter itself. Conversely, values near 1 highlight significant parameter deviations, pointing to more dynamic and potentially insightful areas of the loss landscape. Then, a binarization process is applied to $G_{\text{global}}$ where elements below $T_{\text{p}}$ are set to 0 and those above are set to 1:
$PG_{\text{global},m}=\begin{cases}0&\text{if }G_{\text{global},m}<T_{\text{p}}\\ 1&\text{otherwise}\end{cases}$ (3)
The threshold $T_{\text{p}}$ directly determines the compression ratio of the model by setting the proportion of parameters to be pruned. $T_{\text{p}}$ can be seen has a hard constraint on model size, which is sometimes needed due to devices’ resource constraints. Given their smaller influence, parameters corresponding to 0 are marked for pruning, whereas those marked with 1 are retained, indicating important search directions within the model parameter space. During each FL round, the $K$ participating clients update $PG_{\text{global}}$ by incorporating their $G_{\text{local}}$ deviation values derived from the initial loss exploration phase.
To select an appropriate $T_{\text{p}}$, consider the desired compression ratio for the model. This ratio reflects the extent to which the model needs to be compressed while maintaining acceptable performance. By carefully selecting $T_{\text{p}}$ based on the desired compression ratio and empirical validation, we can achieve a well-balanced model that is both efficient and accurate, tailored to the specific needs of the FL task. In appendix B we conduct ablation studies on $T_{\text{p}}$.
The Proposed AutoFLIP Framework. Here our aim is to argument how the parameter pruning mechanism based on loss exploration enters a general FL edge training framework. Algorithm 1 provides an overview of the entire framework of the proposed AutoFLIP algorithm for FL. It is composed by the following steps.
Algorithm 1 AutoFLIP Algorithm
1:Server Initialization: Initial matrix $W_{\text{global}}^{(0)}$, number of clients for exploration $C_{\text{exp}}$, exploration epochs $E_{\text{exp}}$, pruning threshold $T_{\text{p}}$, FL rounds $R$, training epochs $E$, number of selected clients per round $K$
2:Server selects $C_{\text{exp}}$ clients for exploration
3:$G_{\text{local}_{i}}=(W_{i}^{\text{Initial}}-W_{i}^{\text{Final}})^{2},\forall i\in[1,C_{\text{exp}}]$
4:for round $r=1$ to $R$ do
5:     Server selects $K$ clients
6:     Compute $G_{\text{global}}^{(r)}$ using Eq. (2)
7:     Compute mask $PG_{\text{global}}^{(r)}$ using Eq. (3)
8:     for client $k=1$ to $K$ do
9:         $W_{k,\text{pruned}}^{(r)}=W_{k}^{(r)}\odot PG_{\text{global}}^{(r)}$
10:         for each local epoch $e=0$ to $E-1$ do
11:              $W_{k,\text{pruned}}^{(e+1)}=W_{k,\text{pruned}}^{(e)}-\eta\nabla L_{k}\left(W_{k,\text{pruned}}^{(e)}\right)$
12:         end for
13:     end for
14:     $W_{\text{global}}^{(r)}=\frac{1}{K}\sum_{k=1}^{K}W_{k,\text{pruned}}^{(E)}$triangleright This can be replaced with other FL aggregation algorithms
15:end for
Server initialization (Line 1). The server is initialized with a global model that it is sent to all the clients. At this stage, the total number of clients undergoing exploration, the number of exploration epochs, and the pruning threshold are also decided.
Exploration phase (Lines 2–3).
The preliminary exploration phase aimed at understanding the relevance of each parameter (weight or bias) in view of loss improvement starts. For each client participating (in this study we select all the available clients), a local guidance matrix storing parameter deviations is computed.
Mask update (Lines 5–7). A FL round starts. The server selects $K$ clients that participate in the round. Only the local guidance matrices of those clients are considered to compute a global guidance matrix, which is then used to generate a binary mask for pruning. The mask contains ones only for the parameters with normalized deviations higher than a prescribed threshold $T_{p}$.
Pruning (Lines 8–9). During each round, clients use the pruning mask to compress their models. This happens through element-wise multiplication between their weight matrix and $PG_{\text{global}}$ at that FL round. Parameters aligned with a 0 in $PG_{\text{global}}$ are pruned; those corresponding to a 1 are kept.
FL round with reduced client models (Lines 10–14). The standard algorithm FedAvg [38] is usedin this framework but AutoFLIP can be applied to other SotA FL aggregation algorithm. The pruned clients are trained. The server receives the local model updates and, upon aggregation, proceeds to update the global model with the FL aggregation strategy. Once updated, the global model is either ready for the next communication round or deemed ready for deployment if the convergence criteria are satisfied.
4.2 Robustness and Efficiency of AutoFLIP
Referring to [12, 11, 57, 60], we base our convergence guarantees on a federated stochastic aggregation scheme. The authors’ assumptions on Lipschitz smoothness, convexity of local loss functions, unbiased gradient estimators, finite client answering times, and specific client aggregation weights form the theoretical backbone of AutoFLIP. These conditions ensure that the learning process remains stable and converges efficiently even in the presence of non-IID data distributions. With AutoFLIP, at each round, each client experiences the same pruning strategy with the pruning mask $PG_{\text{global}}$, resulting in a substantial decrease in the variance ( $\sigma^{2}_{\Delta W}$) previously defined in Section 3.2 of weight updates for the global model. This uniform pruning strategy minimizes discrepancies in weight adjustments across clients by focusing updates on critical weights identified during the federated loss exploration phase. The reduction in variance helps to alleviate the bias caused by the non-IID setting, as shown in the work of [65], thus promoting better global convergence.
Figure 2: Distribution of parameter deviations in $G_{global}$ after exploration. Absolute frequency in log-scale is shown for each normalized deviation. Higher frequencies are recorded for smaller deviation values, indicating that many parameters are irrelevant for loss improvement.
Furthermore, [59, 45, 53, 24] provide a theoretical foundation for which pruned NNs can effectively learn signals. They demonstrate that pruning preserves the signal’s magnitude in features and reduces noise, leading to improved generalization. These studies highlight that pruning, when done correctly, does not degrade the model’s capacity to learn but rather focuses the learning on more relevant features. By focusing on parameters with significant contributions to the loss topology, AutoFLIP ensures that the essential features are retained. As illustrated in Figure 2 for different NN, the parameters in $G_{global}$ with minimal variability during the federated loss exploration phase are pruned, while those exhibiting high deviations are retained. Note that higher frequencies are recorded for smaller deviation values, indicating that many parameters, according to our pruning strategy, are non important. The high density of those parameters lead to the weights of entire channels or layers being set to zero. In such cases, we delete the entire substructure of the NN, resulting in a structured network reduction. This leads to an improvement in computing efficiency as detteiled in appendix  E.
AutoFLIP enhances communication efficiency in FL by reducing model sizes transmitted between clients and the server, thus lowering bandwidth requirements for FL rounds. Its selective updating mechanism ensures only essential parameters, those significantly affecting global model convergence, are communicated. The appendices  F,  G and  H provide a detailed analysis of how AutoFLIP accelerates inference and improves training efficiency, demonstrating its significant role in lowering computational costs and boosting FL’s overall applicability.
5 Experiments
Inspired by [16], we benchmark AutoFLIP across established datasets to evaluate its robustness in various non-IID environments. We explore three distinct partitioning approaches for creating strongly non-IID conditions: a Pathological non-IID scenario, which involves clients using data from two distinct classes, employing MNIST with a six-layer CNN (7,628,484 parameters) and CIFAR10 with EfficientNet-B3 (10,838,784 parameters), a Dirichlet-based non-IID scenario, which utilizes the Dirichlet distribution to distribute data among clients, with varying class counts per client, using CIFAR100 with ResNet (23,755,900 parameters), and a LEAF non-IID scenario, which adopts the LEAF benchmark [4] with FEMNIST and Shakespeare datasets. For FEMNIST, a CNN architecture with 13,180,734 parameters is used. For Shakespeare, we consider a two-layer LSTM model with 5,040,000 parameters. Further details on these scenarios are provided in Appendix C.
5.1 Experimental Setup and Results
We evaluate AutoFLIP against both FedAvg without any model compression and with SotA algorithm e.g PruneFL [26], and FL-pruning with various parameter selection criteria: Random, L1, L2, Similarity, and BN mask, as described in [58]. The experimental setup involves $C=20$ (for LEAF non-IID scenario we employ $C=730$ for Shakespeare and $C=660$ for FEMNIST) clients, a batch size $B=350$, a local update epochs $E=5$, and a learning rate $\eta=0.0003$ over $200$ total rounds $R$ with $K=5$ (for LEAF non-IID scenario $K=20$) clients selected per round. We incorporate a server momentum of $0.9$ and use an SGD optimizer with weight decay. The exploration phase consists of up to $E_{\text{exp}}=150$ epochs, and the pruning threshold is set to $T_{\text{p}}=0.3$. Data is divided into 80% for training and 20% for testing, with global model performance assessed by the average prediction accuracy on the test sets. To ensure statistical validity, each experiment is repeated 10 times. We measure the compression rate to evaluate model size reduction and its impact. Experiments were conducted with an Intel Xeon X5680, 128 GB of DDR4 RAM, and an NVIDIA TITAN X GPU.
Figure 3: Average accuracy convergence profiles for the global model within the FL framework.
Pathological non-IID) Here, AutoFLIP achieves an average client compression rate of x1.74. At each round, we remove on average 3244298 parameters of the six-layer CNN for each participant client. For the EfficientNet-B3, we obtain an average compression rate of x2.1 with 5677458 deleted parameters. For a fair comparison with the baselines, we ensure that the number of parameters pruned matches the compression ratio of AutoFLIP, quantified as 42% for the six-layer CNN and 52.38% for EfficientNet-B3.
The first two subplots in Figure 3 show the evolution of global model accuracy during the FL rounds for the four-layer CNN with the MNIST dataset and for EfficientNet-B3 with the CIFAR10 dataset. Refer to Appendix D for the evolution of the loss metric. In the case of the MNIST, the early rounds of FL show that AutoFLIP achieves slightly higher accuracy compared to both FedAvg and the other FL pruning strategies, among which RandomPruning emerges as the top performer. This indicates a faster convergence rate for our proposed method. However, the performance of the baselines soon becomes comparable, with no clear superiority as the FL procedure progresses. We attribute this to the simplicity of the prediction tasks on the MNIST dataset compared to the excessive complexity of the four-layer CNN, which already possesses extremely good prediction capabilities that cannot be further enhanced by pruning. For the CIFAR10 dataset, we do not observe any advantage in using AutoFLIP over the other baselines. Surprisingly, all methods exhibit severe fluctuations in the accuracy convergence profiles up to FL round 100, after which they stabilize and become comparable.
Dirichlet-based non-IID) For ResNet, AutoFLIP achieves an average compression rate for the clients of x1.58, with 8,720,520 parameters pruned on average out of 23,755,900 total parameters. Hence, we adjust the percentage of parameters to be pruned to 36.71% for the different baselines.
The third subplot in Figure 3 illustrates the evolution of the global model accuracy during the FL rounds for ResNet on CIFAR100. Here, AutoFLIP exhibits a performance enhancement throughout the considered training rounds. At round 200, it achieves an accuracy of 0.987, compared to 0.918 for FedAvg and 0.925 for PruneFL. This enhancement signifies the robustness of AutoFLIP, showcasing its ability to maintain elevated performance levels when integrated with larger-complex neural networks and larger datasets.
LEAF non-IID) In this scenario, AutoFLIP achieves an average compression rate of x1.8 for 5858104 client parameters pruned out of 13180734. Hence, we adjust the number of parameters to be pruned for the different baselines to 44%. As observed in the last two subplots of Figure 3 for the FEMNIST and Shakespeare datasets, AutoFLIP consistently outperforms the other pruning strategies by a significant margin.
What stands out is the initial acceleration in convergence speed observed for AutoFLIP, firmly establishing it as a superior choice overFedAvg and the other FL baselines. Furthermore, this superiority persists throughout the entire FL training procedure. The final average accuracy values are 0.985 for AutoFLIP, 0.905 for FedAvg, and 0.935 for RandomPruning on the FEMNIST dataset. For the Shakespeare dataset, the values are 0.815, 0.783, and 0.738, respectively. Here, even L1 proves to be competitive, reaching a final accuracy equal to 0.802. However, it demonstrates inferior initial convergence.
6 Conclusion and Limitations
We introduced AutoFLIP, an innovative federated learning (FL) approach that employs informed pruning to optimize deep learning (DL) models on clients with limited computational resources. Through extensive experiments in various non-IID scenarios, AutoFLIP has demonstrated its ability to achieve better accuracy and significantly reduce computational and communication overheads. It enhances global convergence in federated settings and shows remarkable adaptability and scalability across diverse DL model architectures and multi-class datasets, particularly as the complexity of tasks increases.
Limitations. AutoFLIP shows promise but has limitations. It is primarily tested in the popular efficient single-server setting, not accounting for multi-server or hierarchical environments with diverse client capabilities and model structure. Our tests also assume standard conditions without data label noise.
Future Research Directions. AutoFLIP underscores its potential for future research avenues, such as leveraging loss exploration for guiding complex Neural Architecture Search (NAS) tasks. Enhancements will focus on refining AutoFLIP’s dynamic and adaptive pruning to better client personalization. We aim to perform comparison analysis with other strategies from other domain such us like NAS or Client Dropout. Further, the impact on data privacy and defense against adversarial clients during the federated loss exploration phase has to assessed. Research will also explore the extension of AutoFLIP to more complex DL architectures and its integration into real-world applications across various domains such as healthcare and mobile computing.
Broader Impact. AutoFLIP enhances sustainability and efficiency in FL, reducing the energy footprint of training deep learning models. Its utility in sensitive sectors like healthcare and finance emphasizes its societal importance. However, deploying AutoFLIP requires careful consideration of ethical issues, including data privacy and biases. Proactive management and regulation are crucial to ensure its positive societal impact and responsible integration into critical fields.
Appendix A Ablation Study on $C_{\text{exp}}$ and $E_{\text{exp}}$
We perform an ablation study to assess the sensitivitiy of our method to the parameters $C_{\text{exp}}$ and $E_{\text{exp}}$. The number of explorer clients $C_{exp}$ influences the comprehensiveness of $G_{\text{global}}$ in capturing the intricacies of the clients loss landscapes. The depth of the exploration phase, quantified by the number of exploration epochs $E_{exp}$ influences loss function surface understanding and $G_{\text{global}}$’s knowledge depth.
In particular, we check how the average accuracy and loss for the global model predictions vary for $C_{\text{exp}}\in\{0.25,0.5,0.75,2.0\}$ and for $E_{\text{exp}}\in\{150,300,500,750,1000\}$. We do this for datasets FEMINST the LEAF non-IID scenario 4. It is possible to observe in Figure 4, the x-axis represents the $E_{exp}$ parameter. Each distinct plot corresponds to a different $C_{exp}$, with the highest accuracy achieved distinctly highlighted. In particular, our findings reveal that even a conservative value of $C_{exp}$ can boost the accuracy. The influence of $C_{exp}$ is substantial, with higher counts of explorer clients resulting in improved initial accuracy, indicative of a more robust $G_{\text{global}}$ at the outset of the learning process. A discernible trend suggests that increasing the $E_{exp}$ value generally leads to an improvement in accuracy. However, at higher values, the increase is reduced, indicating a saturation point beyond which additional exploration epochs no longer improve accuracy.
Figure 4: Ablation on the total number of clients for exploration $C_{\text{exp}}$ and the total number of exploration epochs $E_{\text{exp}}$ for FEMNIST/pathological non-IID data, based on average accuracy.
Appendix B Ablation Study on $T_{p}$
We perform an ablation study to assess the sensitivitiy of our method to the pruning threshold parameter $T_{p}$. In particular, we check how the average accuracy and loss for the global model predictions vary for $T_{p}\in\{0.1,0.2,0.3,0.4,0.5\}$. We do this on two datasets: MINST in Figure 5 and CIFAR10 in Figure 6 from the Pathological non-IID scenario. In both cases, $T_{p}=0.3$ seems the most convenient choice.
Figure 5: Ablation on $T_{p}$ for MNIST/non-IID based on average accuracy (top) and loss (bottom).
Figure 6: Ablation on $T_{p}$ for CIFAR10/non-IID based on average accuracy (top) and loss (bottom).
Appendix C Partitioning approaches
Pathological non-IID
This experimental configuration is delineated by each client possessing data exclusively from two distinct classes within a broader multi-class dataset. Figure 7 illustrates this "pathological" data partitioning scenario within the CIFAR10 dataset across 20 clients. For our experiments, we select the MNIST dataset [6] with a six-layer CNN (7628484 parameters) and the CIFAR10 dataset [28] with EfficientNet-B3 architecture (10838784 parameters), following the guidelines in [38] and [52].
Figure 7: Illustration of pathological data partitioning on CIFAR10 for 20 clients, where each color represents a different class.
Dirichlet-based non-IID
This advanced experimental setup, as introduced by [22], utilizes the Dirichlet distribution, modulated by a concentration parameter $\alpha$. Let $\boldsymbol{p}=(p_{1},p_{2},...,p_{N})$ be the class distribution for a given client, where $N$ is the number of classes. The Dirichlet distribution is defined as $\boldsymbol{p}\sim\text{Dir}(\alpha\cdot\boldsymbol{1}_{N})$, where “Dir” denotes the Dirichlet distribution, $\alpha$ is the concentration parameter, and $\boldsymbol{1}_{N}$ is a N-dimensional vector of ones. In this context, a low value of $\alpha$, or $\alpha\to 0$, leads to distributions where most of the probability mass is concentrated on a single class, thereby indicating that each client’s data is restricted to a single class. Conversely, as $\alpha\to\infty$, $\boldsymbol{p}$ approaches a uniform distribution, ensuring that the samples are evenly split across all clients. Figure 8 illustrates this “Dirichlet-based non-IID” data partitioning scenario within the CIFAR100 dataset across 20 clients, with individual colors denoting separate classes.
Figure 8: Illustration of Dirichlet-based non-IID data partitioning on CIFAR100 for 20 clients, where each color represents a different class.
To address the complexities of larger datasets, we have extended our evaluation to include CIFAR100 [28] with a $\alpha=100$, employing ResNet (23755900 parameters) [19] in alignment with the methodology proposed in [16].
LEAF non-IID
Utilizing the popular LEAF benchmark for FL [4], we selected the FEMNIST and Shakespeare datasets to simulate closer real-world FL scenarios, with each dataset designed for specific tasks. The FEMNIST dataset is defined for a multi-class classification challenge involving 62 distinct classes. Conversely, the Shakespeare dataset is tailored for a next-character prediction task, requiring models to predict the subsequent character from a sequence of 80 characters, thereby testing the model capabilities in sequential data processing and language modeling. The incorporation of the next-character prediction task allows for a comprehensive assessment of AutoFLIP adaptability and performance across diverse task types and deep neural network architectures, such as Long Short-Term Memory (LSTM) networks.
In our experimental setup, we employed the FEMNIST-CNN architecture, as delineated in [4], for the FEMNIST dataset. For the Shakespeare dataset, we utilized a two-layer (LSTM) (5040000 parameters) model, in accordance with the specifications provided in [38].
Appendix D Loss plots
We present in Figure 9 the loss convergence profiles for the global model participating in the FL procedure.for the global model. Here, we compare AutoFLIP to the different federated pruning strategies evaluated on both image recognition and text prediction tasks using five distinct datasets: MNIST, CIFAR10, CIFAR100, FEMNIST, and Shakespeare. Due to the varying complexities of each task, we use different model structures for different datasets.
Figure 9: Average loss convergence profiles for the global model within the FL framework.
Appendix E Structured Pruning Experiments
We present an evaluation of AutoFLIP using exclusively structured pruning methods. We compare these results with both the hybrid approach of AutoFLIP and the SotA algorithm PruneFL [26]. We conducted experiments on FEMNIST and CIFAR-100, utilizing the same non-IID settings, model architectures, and parameters as described in Sec. 5.
With structured pruning, we focus solely on removing entire filters, channels, or neurons based on the knowledge derived from the federated loss exploration phase. Unlike the hybrid approach, individual weights within structures were not pruned unless the entire substructure was considered unimportant.
Tables 1 and 2 summarize the performance metrics of AutoFLIP with structured pruning compared to AutoFLIP with hybrid pruning, PruneFL, and the standard FedAvg without model compression.
Table 1: Performance Comparison on FEMNIST with Structured Pruning
Metric AutoFLIP AutoFLIP PruneFL FedAvg
(Structured) (Hybrid) (No Compression)
Compression Rate (%) 33.89
Training Time per Client (s) 15.66 17.39 21.39
Inference Time per Client (ms) 7.5 11.61 14.61
Computation Cost (GFLOPs) 10.08 11.28 19.36
Final Test Accuracy (%) 93.9 98.5 89.3 90.5
Table 2: Performance Comparison on CIFAR-100 with Structured Pruning
Metric AutoFLIP AutoFLIP PruneFL FedAvg
(Structured) (Hybrid) (No Compression)
Compression Rate (%) 49.22
Training Time per Client (s) 186.00 192.01 257.18
Inference Time per Client (ms) 8.1 9.8 15.97
Computation Cost (GFLOPs) 14.39 14.98 17.78
Final Test Accuracy (%) 94.9 98.7 90.6 91.8
The structured pruning approach in AutoFLIP resulted in notable reductions in training time per client. For FEMNIST, the training time decreased by approximately 26.8% compared to FedAvg without compression, and by 10.0% compared to PruneFL. For CIFAR-100, the training time decreased by approximately 27.7% compared to FedAvg, and by 3.1% compared to PruneFL.
Inference times per client were significantly reduced. For FEMNIST, the inference time decreased by 48.7% compared to FedAvg, and by 35.4% compared to PruneFL. For CIFAR-100, the inference time decreased by 49.3% compared to FedAvg, and by 17.3% compared to PruneFL.
The computation cost, measured in giga floating-point operations (GFLOPs), was substantially reduced. For FEMNIST, the computation cost decreased by 47.9% compared to FedAvg, and by 10.6% compared to PruneFL. For CIFAR-100, it decreased by 19.0% compared to FedAvg, and by 4.0% compared to PruneFL.
Regarding memory usage and the number of processed parameters, on FEMNIST, the structured pruning version processed approximately 15.3 billion parameters during the entire federated learning procedure, compared to 28.3 billion for FedAvg. On CIFAR-100, the structured pruning version processed approximately 3.9 billion parameters, compared to 9.4 billion for FedAvg. These reductions contribute to decreased memory bandwidth requirements and potential energy savings, which are significant factors in large-scale federated learning deployments.
The accuracy for FEMNIST was 93.9%, which is lower than that of hybrid AutoFLIP (98.5%), but higher than those of PruneFL (89.3%) and FedAvg (90.5%). For CIFAR-100, the final test accuracy was 94.9%, lower than that of hybrid AutoFLIP (98.7%), but higher than those of PruneFL (90.6%) and FedAvg (91.8%). The hybrid AutoFLIP approach achieves superior accuracy by also eliminating individual weights that may contribute to overfitting or biased learning. Thiss demonstrate that the hybrid pruning approach of AutoFLIP offers the best overall performance by combining the benefits of both structured and unstructured pruning.
Appendix F Inference acceleration
In this section, we discuss the inference acceleration of AutoFLIP. When performing inference on the client’s side with the pruned sub-model, we accelerate the inference time and reduce the computational consumption. Figure 10 shows the inference acceleration comparison after applying AutoFLIP. Notably, the FLOPs (floating point operations per second) in all the evaluated models are reduced. Table 3 shows that the Six-layer CNN deployed for the pathological non-IID experiment with MNIST, experienced a substantial decrease in computational load, equal to a 41.62% reduction in FLOPs. EfficientNet-B3, used for CIFAR10 in the pathological non-IID experiment, saw further improvements, reaching a FLOPs reduction of 46.44%. The deeper ResNet model, designed for CIFAR100 in the Dirichlet-based non-IID experiment, achieved a significant reduction in FLOPs, over 50%, highlighting the potential of AutoFLIP to streamline deep networks for more efficient inference. The FEMNIST-CNN and LSTM models, employed for the LEAF non-IID experiment, showcased a FLOPs reduction equal to 56.49% and 44.44%, respectively.
Figure 10: Original FLOPs and reduced FLOPs
Table 3: FLOPs comparison
Model Compression Rate Original FLOPs Reduced FLOPs FLOPs % Reduced
Six-layer CNN 1.74 13.25 G 5.43 G 41.62% $\downarrow$
EfficientNet-B3 2.1 15.67 G 7.20 G 46.44% $\downarrow$
ResNet 1.58 7.83 G 4.07 G 52.75% $\downarrow$
FEMNIST-CNN 1.8 19.36 G 10.08 G 56.49% $\downarrow$
LSTM 1.8 10.08 G 4.43 G 44.44% $\downarrow$
Appendix G Training efficiency
To ascertain AutoFLIP’s impact on enhancing training efficiency within FL frameworks, we delve into an examination of the associated communication costs. For a practical perspective, the deployed models are trained to achieve a 90% accuracy threshold. As presented in [63], the cost function employed for this evaluation is defined as:
[…]
In Table 4, we observe the effectiveness of AutoFLIP in reducing communication costs across various non-IID scenarios with different models and datasets. Notably, the Six-layer CNN model, used in the MNIST dataset for the Pathological non-IID experiment, demonstrated a significant reduction in communication costs by 41.61%, which underscores AutoFLIP’s effectiveness in simpler architectures. This efficiency extends to more complex architectures, like EfficientNet-B3 and ResNet, employed for the CIFAR10 and CIFAR100 datasets respectively for the Dirichlet-based non-IID experiment, which also saw notable cost reductions of 30.93% and 29.88%. Similarly, the FEMNIST-CNN and LSTM models, used in the LEAF non-IID experiment, exhibited reductions in communication costs by 19.54% and 19.29%, respectively. These results highlight AutoFLIP’s broad applicability and substantial impact on training efficiency across a range of model complexities and dataset types.
Table 4: Comparison of the total communication costs
Model Rounds AutoFLIP Rounds NoAutoFLIP Cost AutoFLIP Cost NoAutoFLIP % Cost Reduced
Six-layer CNN 3 58 189.45 GB 324.43 GB 41.61% $\downarrow$
EfficientNet-B3 27 39 290.26 GB 420.27 GB 30.93% $\downarrow$
ResNet 7 49 712.70 GB 1016.40 GB 29.88% $\downarrow$
FEMNIST-CNN 280 348 369.06 GB 458.69 GB 19.54% $\downarrow$
LSTM 243 301 122.47 GB 151.74 GB 19.29% $\downarrow$
Appendix H Computation Cost
To evaluate AutoFLIP’s role in reducing computational effort, we investigate the number of parameters processed for a single client. Distinguishing between computational efforts on the global model and the clients is essential, with a particular focus on the client side. For AutoFLIP, each client handles a substantial number of parameters over an additional 150 exploration epochs ($E_{\text{exp}}$). From a practical standpoint, we compare AutoFLIP and FedAvg with RandomPruning with the same compression rate. The models are trained to meet a 90% of global accuracy. We define the computation cost function as:
[…]
In Table 5, the pathological non-IID experiment with MNIST using the Six-layer CNN model shows a significant reduction in computational cost by 62.51%. This efficiency extends to more complex architectures like EfficientNet-B3 and ResNet, used for the CIFAR10 and CIFAR100 datasets respectively, with cost reductions of 46.41% and 58.22%. Similarly, the FEMNIST-CNN and LSTM models, employed in the LEAF non-IID experiment, demonstrated reductions in computational costs by 45.99% and 29.60% respectively. These results underline AutoFLIP’s broad applicability and substantial impact on reducing computational efforts across diverse model architectures and dataset types.
Table 5: Comparison of the total computation costs
Model Processed parameters AutoFLIP Processed parameters NoAutoFLIP % Cost Reduced
Six-layer CNN 535,309,170 1,428,653,880 62.51% $\downarrow$
EfficientNet-B3 2,005,175,040 3,740,891,680 46.41% $\downarrow$
ResNet 3,919,723,500 9,378,097,500 58.22% $\downarrow$
FEMNIST-CNN 15,303,653,440 28,338,578,100 45.99% $\downarrow$
LSTM 5,871,600,000 8,335,200,000 29.60% $\downarrow$